OpenStudy (anonymous):

Solve using perfect square factoring patterns. x^2 – 12x + 36 = 4 {–8, –4} {–2, 10} {4, 8} {10, –2}

OpenStudy (gebooors):

You can Express equation (X-6)^2 = 2^2

OpenStudy (jmark):

x^2-12x+36 = 4 group the like term x^2-12x+36 -4 = 0 x^2-12x+32 factories the term of equation x^2-8x-4x+32 = 0 x(x-8)-4(x-8) = 0 (x-8) (x-4) = 0 x-8 = 0, x-4 = 0 x = 8, x =4 the solution is {4,8} ICSE question papers for class 9 - http://icse.edurite.com/icse-question-papers/icse-question-papers-for-class-9.html